OpenStudy (anonymous):

Algebra help; I have nooo clue how to do fractions. Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y A) -12 B)7/4 C) 19/4 D)24

OpenStudy (anonymous):

plug in x and isolate y right?

OpenStudy (anonymous):

so 6(-2) =7 - 4y

hartnn (hartnn):

-12=7-4y -12-7=-4y -19=-4y y=19/4

OpenStudy (anonymous):

First solving for y 6x = 7 – 4y 4y = 7-6x y = (7-6x)/4 Nos substitute x= -2 y = (7-6(-2))/4 y = (7+12)/4 y = 19/4